Jia Inc. has an odd dividend policy

Jia Inc., has an odd dividend policy. The company has just paid a dividend of $4 per share and has announced that it will increase the dividend by $5 per share for each of the next four years, and then never pay another dividend. If you require a return of 12.00% on the company's stock, how much will you pay for a share today?
Answer #1
YearDividend
1(4+5)=9
2(9+5)=14
3(14+5)=19
4(19+5)=24

Hence current price=Future dividends*Present value of discounting factor(rate%,time period)

=9/1.12+14/1.12^2+19/1.12^3+24/1.12^4

=$47.97(Approx)


answered by: aux

> wrong.

Ve Bi Sun, Nov 7, 2021 2:09 AM